# Ada-VE: Training-Free Consistent Video Editing Using Adaptive Motion Prior

### v1

@@ Abstract
Video-to-video synthesis models face significant challenges, such as ensuring consistent character generation across frames, maintaining smooth temporal transitions, and preserving quality during fast motion. The introduction of joint fully cross-frame self-attention mechanisms has improved character consistency, but this comes at the cost of increased computational complexity. This full cross-frame self-attention mechanism also incorporates redundant details and limits the number of frames that can be jointly edited due to its computational cost. Moreover, the lack of frames in cross-frame attention adversely affects temporal consistency and visual quality. To address these limitations, we propose a new adaptive motion-guided cross-frame attention mechanism that drastically reduces complexity while preserving semantic details and temporal consistency. Specifically, we selectively incorporate the moving regions of successive frames in cross-frame attention and sparsely include stationary regions based on optical flow sampling. This technique allows for an increased number of jointly edited frames without additional computational overhead. For longer duration of video editing, existing methods primarily focus on frame interpolation or flow-warping from jointly edited keyframes, which often results in blurry frames or reduced temporal consistency. To improve this, we introduce KV-caching of jointly edited frames and reuse the same KV across all intermediate frames, significantly enhancing both intermediate frame quality and temporal consistency. Overall, our motion-sampling method enables the use of around three times more keyframes than existing joint editing methods while maintaining superior prediction quality. Extensive qualitative and quantitative evaluations demonstrate the superiority of our adaptive video editing techniques over existing methods. For example, Ada-VE achieves up to 4$\times$ speed-up when using fully-extended self-attention across 40 frames for joint editing, without compromising visual quality or temporal consistency.
@@ 1 Introduction
The field of video generation has gained significant attention due to its numerous practical applications, including video editing [20], video synthesis [35], and style transfer [34]. This paper focuses on video-to-video synthesis, which involves altering videos through text prompts while preserving their original motion and structure [18, 11, 44]. Recently, video synthesis techniques based on diffusion models have garnered considerable interest for their outstanding performance in image editing and generation tasks [38, 3]. However, training these models from scratch is computationally intensive, data-demanding, and necessitates frequent updates with the release of new models [21]. To address these challenges, we explore training-free extension of text-to-image diffusion models for video-to-video synthesis tasks, avoiding the need for extensive retraining and leveraging the capabilities of pre-trained models.
Figure 1: We examine the impact of cross-frame self-attention in video editing under two scenarios: slow motion, where frames are sampled at an interval of 1, and fast motion, where frames are sampled at an interval of 10, representing higher motion variance. Joint editing is performed in both scenarios using fully-extended self-attention KVs and self-attention KVs from only the first frame. Using only the first frame’s KV significantly deteriorates performance in fast motion, while achieving reasonable performance at slow motion. In contrast, fully-extended self-attention performs consistently better in both scenarios but significantly increases computational burden.
In general, video editing faces several key challenges that hinder its widespread adoption for real-world applications. First, consistent character generation across frames based on text prompts, while preserving the structural details of the guidance video, is crucial. Since diffusion models generate outputs stochastically from noise, achieving consistent character generation over long-duration videos (more than $100$ frames) is challenging, often resulting in flickering for non-persistent content across successive frames. Second, maintaining temporal coherence in the generated frames is essential for realism, as human eyes are highly sensitive to temporal inconsistencies. Most existing video diffusion methods suffer from blurring effects [11] and temporal inconsistencies [44, 18], compromising the quality of the video. Lastly, real-world videos exhibit diverse dynamics, including fast motion transitions, necessitating adaptive model operations to maintain quality. A single approach may not be suitable for all video classes in real-world applications, highlighting the need for adaptable and robust methods across different scenarios.
@@
For generating longer-duration videos, most recent works have focused on two-step approaches: first, jointly editing sampled key reference frames, followed by custom intermediate frame-interpolation techniques. Several interpolation methods have been explored, including flow-warping [21], token flow [11], and traditional video interpolation techniques [42]. However, these methods often lead to blurring effects, especially in videos with fast motion, which reduces the overall video quality. Another approach involves extracting relevant features from nearby key frames only during the joint editing phase and selectively reusing these features for intermediate frames [44]. Despite this, such methods typically suffer from inconsistent character generation and temporal inconsistencies, heavily dependent on the feature selection method used. Therefore, the trade-off between motion blurring and temporal inconsistency necessitates more sophisticated methods for longer video generation. Generally, extensive use of cross-frame self-attention can improve performance, but at the cost of increased computational demands, while consistent frame interpolation remains a significant bottleneck for achieving high-quality longer videos.
In this paper, we propose Adaptive Video-Editing (Ada-VE), a novel approach to address the major challenges in video-to-video synthesis. Ada-VE adaptively integrates essential details in joint cross-frame attention while suppressing redundant parts, thereby increasing the number of frames that can be edited without compromising visual quality or increasing computational burden. To achieve this, we selectively incorporate the most salient key-value pairs (KVs) from the moving regions across frames, while leveraging optical flow estimation to identify these regions. This dynamic extension of KVs in cross-frame attention optimizes self-attention based on the motion dynamics of the video, thus reducing redundant computations and enabling faster processing without losing detail. Additionally, we find that preserving the same KVs across frames is crucial for maintaining inter-frame consistency. Based on this observation, we introduce the KV-cache technique, which reuses KVs from cross-frame self-attention in intermediate frames. This method enhances the generation of longer-duration videos, improving both visual quality and temporal consistency.
@@
    We propose an adaptive extension of cross-frame self-attention, enabling processing an increased number of jointly-edited frames without increasing the computational complexity.
@@
    We introduce the use of KV-cache from jointly edited frames to maintain temporal consistency and visual quality in intermediate frames of longer duration videos.
@@
    We present extensive experimental studies that demonstrate the effectiveness of AdaVE. For example, when incorporating Ada-VE on fully-extended self-attention for 40 frames, operational latency is reduced by 70% on average, while maintaining the same warp-error and CLIP-score.
@@ 2.1 Video Synthesis Methods
Many approaches have attempted to inflate cross-frame self-attention to utilize image priors for video synthesis [41, 11, 44]. However, finding the optimal balance between the number of frames and computational efficiency remains a significant challenge in existing methods. More recently, SORA has demonstrated surprising results in video synthesis, but it is not available as an open-source tool and requires large-scale training on curated datasets [23]. Despite these advancements, open-source methods often produce shorter videos (less than 100 frames long), suffer from flickering due to inconsistent character generation, and exhibit reduced visual quality. Our goal is to address these limitations and enhance the performance of video synthesis methods.
2.2 Frame Interpolation and Temporal Consistency
Maintaining temporal consistency and preserving the structure of the original video in the generated frames is a primary challenge in existing methods [20, 19, 44]. Several approaches have introduced flow-warping techniques to generate intermediate frames from jointly edited frames [42, 21, 18]. However, flow-warping often introduces errors due to occlusions. Some methods have attempted to train diffusion models as inpainting models to address these limitations, but consistent inpainting remains challenging and requires large-scale training, often resulting in lower visual quality and flickering [21]. Instead of using flow-warping for interpolation as prior work, we primarily use optical-flow [37, 16, 17] based motion prior to estimate the moving parts, and enhance cross-frame attention performance.
Tune-a-Video [41] proposed fine-tuning the model on a specific video, but this approach sacrifices generalizability. Recently, TokenFlow [11] introduced a method of flowing self-attention features from key frames based on spatial similarity. Despite its promising performance, TokenFlow often blurs intermediate frames and reduces overall quality. ControlVideo [44] proposed reusing attention features from nearby frames to improve consistency. However, this method can still result in significant flickering and inconsistent character generation for feature disparity in subsequent frame synthesis. Our method focuses on enhancing both consistency and quality by optimizing the reuse of self-attention features from joint-editing, aiming to overcome the limitations of previous approaches.
@@ 3 Preliminaries
Figure 2: Ada-VE Overview: (i) During preprocessing, DDIM inversion is performed to extract deterministic noise $\mathbf{X}_{T}\sim\mathcal{N}(0,I)$, and successive motion masks $\mathcal{M}$ are extracted using an optical flow model. (ii) Several reference frames $\mathbf{X}_{t,ref}$ are then sampled at timestep $t$, and jointly edited iteratively with the proposed sparse extension of self-attention KVs guided by motion masks $\mathbf{\mathbf{\mathcal{M}}}$, with all extended KVs being cached. (iii) Finally, all intermediate frames $\mathbf{X}_{t,int}$ are edited using the cached sparse reference KVs at timestep $t$.
@@ 3.1 Diffusion Probabilistic Models (DPM)
DPMs [27, 14, 8, 28] introduce a family of generative models that learn a data distribution through iterative progressive denoising of a Gaussian i.i.d noisy image $\mathbf{X}_{T}\sim\mathcal{N}(0,I)$. DPM generates a clean image ${x}_{0}$ of the target conditional distribution $q$ using additional text guidance $\mathcal{T}$. Subsequent work [33, 3] introduced Latent Diffusion Models (LDMs) to generate high-resolution images by processing the conditional denoising in the latent space. There are multiple components in the LDM architecture, relying on the base U-Net architecture. Initially, a pre-trained encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are trained to extract intermediate latent representations and generate representative images from the latents, respectively. The core denoising components consist of: a residual convolutional block to preserve spatial features, a self-attention (SA) block for feature recalibration, followed by a cross-attention block to introduce text guidance. In particular, SA blocks [39] measure the feature affinity across projected $d$-dimensional query and key features $Q$ and $K$, respectively, followed by recalibration with value features $V$ using:
$\mathtt{SelfAttention}(X)=\mathtt{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (1)
The self-attention plays a critical role in preserving the structures of the generated frames [38, 44, 11].
@@ 3.2 Controlled Editing with LDM
These zero-shot image editing techniques can be applied directly to videos. However, a major challenge is achieving consistent generation across frames. Prior work introduced extended cross-frame self-attention  [11, 30] to preserve generated content by extending the keys and values of all $N$ frames in each self-attention using:
$\mathtt{ExtendedSelfAttention}(X_{i})=\mathtt{Softmax}\left(\frac{Q_{i}K_{\mathtt{all}}^{T}}{\sqrt{d}}\right)V_{\mathtt{all}},\ \ \ \forall i=\{1,2,\dots,N\}$ (2)
$\text{where}\ K_{\mathtt{all}}=\{K_{1},K_{2},\dots,K_{N}\},\ V_{\mathtt{all}}=\{V_{1},V_{2},\dots,V_{N}\}$
While such extensions play a crucial role in consistent character generation, naively extending keys $K_{\mathtt{all}}$ and values $V_{\mathtt{all}}$ consumes significant memory and computational resources. Furthermore, despite achieving consistent character generation, other challenges include generating smooth, temporally consistent frames over the video.
@@ 4 Methodology
To increase the number of frames in joint editing and reduce computational burden, we propose a motion-adaptive KV selection method. This method derives the KV extension directly from the motion guidance of the given video. Additionally, to extend the temporal window for generating longer videos, we introduce the KV-cache technique, which reuses KVs from key pivotal frames during joint editing. Our method builds on Plug-and-Play (PnP) [38] image editing methods, which start with DDIM-inversion of the guidance video, followed by injecting self-attention and residual features during final editing. Our primary contribution is the motion-adaptive extension of the cross-frame self-attention method and the KV-caching technique for enhanced video synthesis (see Fig. 2). The following sections present the proposed approach in detail.
Figure 3: Motion Mask Extraction: An off-the-shelf model is used to extract optical flow from each pair of successive frames. The extracted flow maps are then converted to RGB and subsequently to grayscale images. Finally, a simple thresholding method is applied to extract the moving region masks in each frame.
@@ 4.1 Adaptive Cross-Frame Attention
Since in general video contains redundant information, we selectively perform KV-extension in self-attention based on temporal motion. Moving regions in each frame contain salient information about dynamic structures, whereas stationary regions generally contain repetitive features. Our primary goal is to incorporate more details from the moving regions across multiple frames while sparsely integrating stationary features. This process can be divided into two parts: extracting motion masks from the guidance video using off-the-shelf optical flow models and integrating these masks to selectively extend reference frame KVs in each self-attention block.
Figure 4: (i) Basic Self-Attention: Queries (Q), keys (K), and values (V) of each frame are independently used. (ii) Fully Extended Self-Attention: All keys and values are combined into $K_{all}$ and $V_{all}$ for cross-frame self-attention. (iii) Proposed Sparsely Extended Self-Attention: Keys and values of all frames are sparsely extended into $K_{sparse}$ and $V_{sparse}$ to incorporate more details of moving regions by leveraging motion masks $\mathbf{\mathcal{M}}$.

### v2

@@ Abstract
Video-to-video synthesis poses significant challenges in maintaining character consistency, smooth temporal transitions, and preserving visual quality during fast motion. While recent fully cross-frame self-attention mechanisms have improved character consistency across multiple frames, they come with high computational costs and often include redundant operations, especially for videos with higher frame rates. To address these inefficiencies, we propose an adaptive motion-guided cross-frame attention mechanism that selectively reduces redundant computations. This enables a greater number of cross-frame attentions over more frames within the same computational budget, thereby enhancing both video quality and temporal coherence. Our method leverages optical flow to focus on moving regions while sparsely attending to stationary areas, allowing for the joint editing of more frames without increasing computational demands. Traditional frame interpolation techniques struggle with motion blur and flickering in intermediate frames, which compromises visual fidelity. To mitigate this, we introduce KV-caching for jointly edited frames, reusing keys and values across intermediate frames to preserve visual quality and maintain temporal consistency throughout the video. With our adaptive cross-frame self-attention approach, we achieve a threefold increase in the number of keyframes processed compared to existing methods, all within the same computational budget as fully cross-frame attention baselines. This results in significant improvements in prediction accuracy and temporal consistency, outperforming state-of-the-art approaches. Code will be made publicly available at https://github.com/tanvir-utexas/AdaVE/tree/main.
@@ 1 Introduction
The field of video generation has gained significant attention due to its numerous practical applications, including video editing [21], video synthesis [37], and style transfer [36]. This paper focuses on video-to-video synthesis, which involves altering videos through text prompts while preserving their original motion and structure [18, 11, 47]. Recently, video synthesis techniques based on diffusion models have garnered considerable interest for their outstanding performance in image editing and generation tasks [40, 3]. However, training these models from scratch is computationally intensive, data-demanding, and necessitates frequent updates with the release of new models [23]. To address these challenges, we explore training-free extension of text-to-image (T2I) diffusion models for video-to-video synthesis tasks, avoiding the need for extensive retraining and leveraging the capabilities of pre-trained models.
Figure 1: Effect of self-attention extension on diverse motion: Frames are sampled at intervals of 1 (slow motion) and 10 (fast motion). Two methods are compared: one uses the first frame’s key-value (KV) across all frames for efficiency, while the other fully extends KVs, which is computationally intensive. The efficient method works well in slow motion but struggles with faster motion, where full extension achieves better results. This highlights the need for adaptive self-attention based on motion to enhance video quality, reduce redundant computations, and incorporate more frames in the self-attention process.
In contrast to image editing, video editing faces several key challenges that hinder its widespread adoption for real-world applications. First, consistent character generation across frames based on text prompts, while preserving the structural details of the guidance video, is crucial. Since diffusion models generate outputs stochastically from noise, achieving consistent character generation over the video is challenging, often resulting in flickering for non-persistent content across successive frames. Second, maintaining temporal coherence in the generated frames is essential for realism, as human eyes are highly sensitive to temporal inconsistencies. Most existing video diffusion methods suffer from blurring effects [11] and temporal inconsistencies [47, 18], compromising the quality of the video. Lastly, real-world videos exhibit diverse dynamics, including fast motion transitions, necessitating adaptive model operations to maintain quality. A single approach may not be suitable for all video classes in real-world applications, highlighting the need for adaptable and robust methods across different scenarios.
@@
In addition to consistent editing across frames, the video editing objective has additional challenges of achieving temporal/motion consistency in all intermediate frames. Most recent works have focused on two-step approaches for maintaining temporal consistency across the video: first, jointly editing a set of randomly/uniformly sampled key frames over the video, followed by custom intermediate frame-interpolation techniques. Several interpolation methods have been explored, including flow-warping [23], token flow [11], and traditional video interpolation techniques [44]. However, these methods often lead to either blurring or flickering effects, especially in videos with fast motion, which reduces the overall video quality. Another approach involves extracting relevant features from nearby key frames only during the joint editing phase and selectively reusing these features for intermediate frames [47]. Despite this, such methods typically suffer from inconsistent character generation and temporal inconsistencies, heavily dependent on the feature selection method used. Therefore, the trade-off between motion blurring and temporal inconsistency necessitates more sophisticated methods for robust video editing. Generally, extensive use of cross-frame self-attention can improve performance, but at the cost of increased computational demands, while consistent frame interpolation remains a significant bottleneck for achieving high-quality videos.
In this paper, we propose Adaptive Video-Editing (Ada-VE), a novel approach to address the major challenges in video-to-video synthesis. Ada-VE adaptively integrates essential details in joint cross-frame attention while suppressing redundant parts, thereby increasing the number of frames that can be edited without compromising visual quality or increasing computational burden. To achieve this, we selectively incorporate the most salient key-value pairs (KVs) from the moving regions across frames, while leveraging optical flow estimation to identify these regions. This dynamic extension of KVs in cross-frame attention optimizes self-attention based on the motion dynamics of the video, thus reducing redundant computations without losing details. Additionally, we find that preserving the same KVs across frames is crucial for maintaining inter-frame consistency. Based on this observation, we introduce the KV-cache technique, which reuses KVs from cross-frame self-attention in intermediate frames. Both of these methods contributes improving both visual quality and temporal consistency.
@@
    We propose an adaptive extension of cross-frame self-attention, enabling processing an increased number of jointly-edited frames for better consistency without increasing the computational complexity.
@@
    We introduce the use of KV-cache from jointly edited frames to maintain temporal consistency and visual quality in intermediate frames of the video.
@@
    We conduct extensive experiments demonstrating the effectiveness of Ada-VE compared to state-of-the-art methods. Notably, with the same computational budget, Ada-VE processes three times more frames in extended cross-frame self-attention, leading to a 45% reduction in warp error and a 25% improvement in CLIP-Score.
@@ 2.1 Video Synthesis Methods
Many approaches have attempted to inflate cross-frame self-attention to utilize image priors for video synthesis [43, 11, 47]. However, finding the optimal balance between the number of frames and computational efficiency remains a significant challenge in existing methods. More recently, SORA has demonstrated surprising results in video synthesis, but it is not available as an open-source tool and requires large-scale training on curated datasets [25]. Despite these advancements, open-source methods often suffer from flickering and blurring effects due to inconsistent character generation, and exhibit reduced visual quality. Our goal is to address these limitations and enhance the performance of video synthesis methods.
2.2 Temporal Consistency in Video Editing
Maintaining temporal consistency and preserving the structure of the original video in the generated frames is a primary challenge in existing methods [21, 19, 47]. For consistent character generation across multiple frames, diverse heuristic self-attention extension method is introduced [23, 11, 18]. Several approaches have introduced flow-warping techniques to generate intermediate frames from jointly edited frames [44, 23, 18]. However, flow-warping often introduces errors due to occlusions. Some methods have attempted to train diffusion models as inpainting models to address these limitations, but consistent inpainting remains challenging and requires large-scale training, often resulting in lower visual quality and flickering [23]. Instead of using flow-warping for interpolation as prior work, we primarily use optical-flow [39, 16, 17] based motion prior to coarsely estimate the moving parts to enhance performance, without being constrained by precise optical flow estimation.
Tune-a-Video [43] proposed fine-tuning the model on a specific video, but this approach sacrifices generalizability. Recently, TokenFlow [11] introduced a method of flowing self-attention features from key frames based on spatial similarity. Despite its promising performance, TokenFlow often blurs intermediate frames and reduces overall quality. ControlVideo [47] proposed reusing attention features from nearby frames to improve consistency. However, this method can still result in significant flickering and inconsistent character generation for feature disparity in subsequent frame synthesis. More recently, Motion-Transfer [45] introduced use of pre-trained text-to-video models for editing, requiring additional training on video data. Moreover, this approach significantly increases computational overhead compared to T2I-based methods, restricting its use to very short video duration. Several real-time video editing methods [22, 20] have been introduced, focusing primarily on faster processing with minimal self-attention extension based on predefined heuristics, but these often suffer from flickering and temporal inconsistencies. Our method focuses on enhancing both consistency and quality by optimizing the adaptive-extension and reuse of self-attention features from joint-editing, aiming to overcome the limitations of previous approaches.
@@ 3 Preliminaries
Figure 2: Ada-VE Overview: (i) During preprocessing, DDIM inversion is performed to extract deterministic noise $\mathbf{X}_{T}\sim\mathcal{N}(0,I)$, and successive coarse motion masks $\mathcal{M}$ are extracted using a lightweight optical flow model, in a single step. (ii) Several reference frames $\mathbf{X}_{t,ref}$ are then sampled at timestep $t$, and jointly edited iteratively with the proposed sparse extension of self-attention KVs guided by motion masks $\mathbf{\mathbf{\mathcal{M}}}$, with all extended KVs being cached. (iii) Finally, all intermediate frames $\mathbf{X}_{t,int}$ are edited using the cached sparse reference KVs at timestep $t$.
@@ 3.1 Diffusion Probabilistic Models (DPM)
DPMs [29, 14, 8, 30] introduce a family of generative models that learn a data distribution through iterative progressive denoising of a Gaussian i.i.d noisy image $\mathbf{X}_{T}\sim\mathcal{N}(0,I)$. DPM generates a clean image ${x}_{0}$ of the target conditional distribution $q$ using additional text guidance $\mathcal{T}$. Subsequent work [35, 3] introduced Latent Diffusion Models (LDMs) to generate high-resolution images by processing the conditional denoising in the latent space. There are multiple components in the LDM architecture, relying on the base U-Net architecture. Initially, a pre-trained encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are trained to extract intermediate latent representations and generate representative images from the latents, respectively. The core denoising components consist of: a residual convolutional block to preserve spatial features, a self-attention (SA) block for feature recalibration, followed by a cross-attention block to introduce text guidance. In particular, SA blocks [41] measure the feature affinity across projected $d$-dimensional query and key features $Q$ and $K$, respectively, followed by recalibration with value features $V$. The self-attention plays a critical role in preserving the structures of the generated frames [40, 47, 11].
@@ 3.2 Controlled Editing with LDM
These zero-shot image editing techniques can be applied directly to videos. However, a major challenge is achieving consistent generation across frames. Prior work introduced extended cross-frame self-attention  [11, 32] to preserve generated content by extending the keys and values of each $i^{th}$ frame of total $N$ frames in each self-attention using:
$\mathtt{ExtendedSelfAttn}(X_{i})=\mathtt{Softmax}\left(\frac{Q_{i}K_{\mathtt{all}}^{T}}{\sqrt{d}}\right)V_{\mathtt{all}},$ (1)
$\text{where}\ K_{\mathtt{all}}=\{K_{n}\}_{n=1}^{N},V_{\mathtt{all}}=\{V_{n}\}_{n=1}^{N}$
While such extensions play a crucial role in consistent character generation, naively extending keys $K_{\mathtt{all}}$ and values $V_{\mathtt{all}}$ across all frames consumes significant memory and computational resources. Moreover, despite achieving consistent character generation, other challenges include generating smooth, temporally consistent frames over the video.
@@ 4 Methodology
To increase the number of frames in joint editing and reduce computational burden, we propose a motion-adaptive KV selection method. This method derives the KV extension directly from the motion guidance of the given video. Additionally, to generate temporally consistent intermediate frames, we introduce the KV-cache technique, which reuses KVs from key pivotal frames during joint editing. Our method builds on Plug-and-Play (PnP) [40] image editing methods, which start with DDIM-inversion of the guidance video, followed by injecting self-attention and residual features during final editing. Our primary contribution is the motion-adaptive extension of the cross-frame self-attention method, and the KV-caching technique for enhanced video synthesis (see Fig. 2). The following sections present the proposed approach in detail.
Figure 3: Motion Mask Extraction: A lightweight, off-the-shelf model is employed to extract coarse optical flow maps from each pair of successive frames. These flow maps are converted to RGB and then to grayscale images. Finally, a thresholding technique is applied to extract moving region masks for each frame. This operation is performed once during preprocessing in a single step.
@@ 4.1 Adaptive Cross-Frame Attention
Since in general video contains redundant information, we selectively perform KV-extension in self-attention based on temporal motion. Moving regions in each frame contain salient information about dynamic structures, whereas stationary regions generally contain repetitive features. Our primary goal is to capture more details from moving regions across multiple frames while sparsely integrating stationary features, thereby reducing the reliance on precise motion estimation. This process can be divided into two parts: extract motion masks from the guidance video using off-the-shelf optical flow models, and integrate these masks to selectively extend reference frame KVs in self-attention blocks.
We note that motion mask extraction is performed using real-time, lightweight optical flow models to coarsely highlight moving regions. This extraction is only done once, while the T2I diffusion model runs iteratively over multiple timesteps (e.g., DDIM runs 50 steps), making the computational cost of flow estimation negligible by comparison. Additionally, our sparse integration of stationary regions reduces the need for precise estimation of moving regions with optical flow models, which is often impractical due to challenges such as occlusions, motion blur, and complex background movements.
Figure 4: (i) Basic Self-Attention: Queries (Q), keys (K), and values (V) are independently used for each frame. (ii) Fully Extended Self-Attention: All keys and values are combined into $K_{all}$ and $V_{all}$ for cross-frame self-attention. (iii) Proposed Sparsely Extended Self-Attention: Keys and values from all frames are sparsely extended into $K_{sparse}$ and $V_{sparse}$ to capture more details of moving regions than stationary background regions, utilizing motion masks $\mathbf{\mathcal{M}}$.